40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 22 18:42:10 2011, kiwamu, Summary, Green Locking, Some issues on the Y end green PDH locking 
    Reply  Wed Nov 23 00:00:53 2011, Zach, Summary, Green Locking, Some issues on the Y end green PDH locking 
    Reply  Sat Nov 26 19:07:40 2011, kiwamu, Summary, Green Locking, AM trnasfer function of the Y end laser with doublin crystal AMTF_lightwave.pngAMTF_cailbrated.bod.zip
Message ID: 6018     Entry time: Sat Nov 26 19:07:40 2011     In reply to: 5980
Author: kiwamu 
Type: Summary 
Category: Green Locking 
Subject: AM trnasfer function of the Y end laser with doublin crystal 

Quote from #5980

 As a part of the ALS noise budgeting we took a look at the Y end PDH setup to see if we are limited by an effect from the RF Amplitude Modulation (AM).

The AM transfer function of the Y end laser has been measured again, but using the frequency-doubled laser this time.

Here is the latest plot of the AM transfer function. The Y-axis is calibrated to RIN (Relative Intensity Noise) / V.

IFBW (which corresponds to a frequency resolution) was set to 100 Hz and the data was averaged about 40 times in a frequency range of 100 kHz - 400 kHz.

Also the zipped data is attached.

AMTF_lightwave.png

It is obvious that out current modulation frequency of 179 kHz (178850 Hz) is not at any of the notches.

It could potentially introduce some amount of the offset to the PDH signal, which allows the audio frequency AM noise to couple into the PDH signal.

Currently I am measuring how much offset we have had because of the mismatched modulation frequency and how much the offset can be reduced by tuning the modulation frequency.

Attachment 2: AMTF_cailbrated.bod.zip  11 kB
ELOG V3.1.3-